If Mary can paint a house in 4 hours and Matt in 6 hours, how long will it take them to paint the house together? How would you set up the equation for this problem?
Answer by richwmiller(17219)

How would you set up the equation for this problem?
Correctly.
x/4 +x/6=1
3x+2x=12
5x=12
x=12/5=2 2/5= 2.4 hrs.